# PASER [Uncaptioned image]: Post-Training Data Selection for Efficient Pruned Large Language Model Recovery

### v1

@@ 3.4 Negative Transfer Mitigation
A CCG is a graph $G=(V,E)$ where vertices $V$ represent concepts, and an edge $(v_{i},v_{j})\in E$ exists if concepts $v_{i}$ and $v_{j}$ co-occur in at least one instruction tuning sample without conflict.
For each new sample $(x,y)$, we extract its concept $C(x,y)$ and check for consistency: $\text{IsConsistent}(x,y)=\forall v_{i},v_{j}\in C(x,y):(v_{i},v_{j})\in E\text{ or }\{v_{i},v_{j}\}\not\subset V$. We only add samples that are consistent with the existing CCG, updating the graph with each addition. This approach ensures that we maintain a coherent set of instructions, minimizing the risk of negative transfer by avoiding the introduction of conflicting concepts during the recovery training process. The full version of our algorithm is provided in Appendix B.
@@ 3.5 Time Complexity Analysis
We provide a comprehensive analysis of the time complexity for the PASER algorithm.
@@
Proof. We analyze each step of the algorithm in detail: The Semantic-Structural Recovery Instruction Clustering step, including SentenceBERT embedding, Diffusion Kernel computation, and NMF Spectral Clustering, has a dominant complexity of $O(N\log N)$. For the Capability Degradation Assessment step, computing JSD for each sample and calculating CDS for each cluster take $O(N)$ time in total. The Inter-capability Budget Allocation, which involves allocating the budget to clusters, has a time complexity of $O(K)$, where $K$ is the number of clusters. However, since $K\leq N$, this step does not dominate the overall complexity. During Intra-capability Efficiency-Driven Sample Selection, for each cluster $c_{k}$, we perform sorting by JSD ($O(|c_{k}|\log|c_{k}|)$), iterate through sorted samples ($O(|c_{k}|)$), perform consistency checks (IsConsistent, $O(C^{2})$ per sample), and update the CCG ($O(C^{2})$ per sample). Considering all clusters, this step’s total complexity is $O(N\log N+NC^{2})$. Thus, the overall time complexity is dominated by the clustering step and the intra-capability sample selection step. Therefore, the total time complexity is $O(N\log N+NC^{2})$.
In practice, $C$ is often much smaller than $N$ ($C\ll N$) and can be considered as a constant factor for large $N$. Thus, we can simplify the complexity to $O(N\log N)$. This analysis demonstrates that PASER is computationally efficient and scalable for large instruction tuning datasets.
@@ 4.1 Experiment Setup
Target LLMs The experiments are performed on several open-source popular English LLMs: LLaMA2-7B/13B/70B (Touvron et al., 2023) (hf version), LLaMA3-8B (Dubey et al., 2024)(instruct version), and bilingual LLMs: Baichuan2-7B/13B (Yang et al., 2023)(base version), which support both English and Chinese.
@@
Base Pruning Schemes Different pruning schemes are incorporated into our experiments to explore the applicability of PASER, ranging from structured pruning methods: LLM-Pruner (Ma et al., 2023), SliceGPT (Ashkboos et al., 2024), semi-structured pruning method: Wanda (Sun et al., 2024), and unstructured pruning method: SparseGPT (Frantar & Alistarh, 2023).
@@
Evaluation Datasets and Tasks To thoroughly evaluate the performance of recovered LLMs, we employ seven common sense reasoning datasets:BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018). In the practice, we relies on the open-source library [^1] https://github.com/EleutherAI/lm-evaluation-harness to implement the evaluation, where the model needs to rank the choices in the multiple choice tasks or generate the answer in the open-ended generation tasks. The whole process is conducted in the zero-shot manner. Besides, we follow  (Ma et al., 2023) to evaluate the language modeling capability with the zero-shot perplexity (PPL) analysis on WikiText2 (Merity et al., 2022) and PTB (Marcus et al., 1993). For enhancing evaluation comprehensiveness, we also conduct experiments on mathematical problem solving and code generation tasks, as shown in Appendix C and D, respectively.
@@ 4.2 Experiment Results and Analysis
Recovery Performance for Different Pruning Schemes We evaluate the recovery performance of LLaMA2-7B using different instruction tuning data selection methods under structured pruning, semi-structured pruning, and unstructured pruning, respectively. According to the results in Table 1, directly employing full data can indeed bring the sub-optimal recovery performance, especially under the LLM-Pruner. This is because the full version of data contains some irrelevant or conflicting information for capability recovery, resulting in the negative transfer during the training phase. Meanwhile, even the general instruction tuning data selection methods like IFD and Nuggets can bring better reasoning and language model performance than full data and random in most cases, validating the necessity of conducting recovery data selection. Furthermore, we can find that previous selection methods can hardly help model recover to the level of unpruned status, under the limited data budget. However, our PASER can not only outperform baselines, but also reduce the averaged reasoning performance degradation to less than 3% under LLM-Pruner, Wanda, and SparseGPT. Especially, when pruning LLaMA2-7B with SliceGPT, our PASER can improve the average reasoning performance to 64.31, higher than the unpruned model. Besides, its zero-shot perplexity on WikiText2 and PTB is also lower than unpruned model slightly. This suggests that allocating recovery budget according to capability degradation levels and prioritizing most-affected samples exhibit the potential of help pruned LLM recover to the capability level of unpruned status. Besides, PASER can also be extended to other LLM post-compression scenarios, like the post-quantization recovery. The corresponding results and analysis are provided in Appendix E.
@@
Robustness over Various Target Large Language Models To validate whether PASER can maintain the robust effectiveness among various target LLMs, we conduct the experiments on LLaMA2-7B/13B/70B, LLaMA3-8B, Baichuan2-7B/13B, under LLM-Pruner. According to results in Table 2, we can first observe the model capability under high pruning ratio (50%) is hard to recover to unpruned level, especially for relatively smaller model like LLaMA2-13B and Baichuan2-13B. Though, PASER can still outperform random and best-performing data selection baseline, Nuggets, by 4.41 and 2.31 points, respectively on average. Especially, for LLaMA2-70B, our PASER can control the averaged reasoning performance degradation to less than 3%. This can be explained that the structure redundancy in 70B model is relatively higher, paving the way for effective recovery through data selection under high pruning ratios. As for the second smallest model, LLaMA3-8B, PASER can recover the reasoning performance to the 99.56% of the unpruned status, which further demonstrates the robustness of PASER over different target LLMs. Finally, the performance of various recovery methods including PASER on Baichuan2-7B is not satisfying enough given only 25% pruning ratio, which can be attributed to that the pruning process has severely damaged the model internal structure.
Table 3: Recovery performance of different instruction tuning data selection methods under two structured pruning schemes on LLaMA2-7B model. The ‘bold’ represents the best performance under the same pruning scheme. Here, the LaMini is taken as the original dataset. The “Reason” indicates the averaged performance on 7 reasoning datasets
@@
Figure 2: Average reasoning performance and recovery post-training time consumption curves corresponding to different instruction tuning data selection methods. The left two subfigures are for Alpaca while right two are subfigures for LaMini.
@@
Ablation Study To validate the contribution of each component in PASER, we conduct comprehensive ablation studies. Specifically, we evaluate three variants: (1) PASER w/o S2RIC: replacing semantic-structural clustering with random clustering while keeping other modules unchanged; (2) PASER w/o CDAIS: randomly sampling equal number of instructions from each cluster instead of using capability degradation-aware selection; (3) PASER w/o NTM: removing the negative transfer mitigation module. The results in Table 4 demonstrate that all three components contribute positively to model recovery across different pruning schemes. The semantic-structural clustering effectively identifies capability-specific instruction groups, leading to 0.18-1.43 points improvement in reasoning performance. Its removal causes degradation in both language modeling (increased perplexity) and reasoning tasks, particularly evident under structured pruning schemes like LLM-Pruner and SliceGPT. The capability degradation-aware selection mechanism enhances recovery efficiency through adaptive budget allocation, contributing 0.29-1.63 points improvement in reasoning tasks while maintaining stable language modeling performance. Notably, negative transfer mitigation shows significant impact (0.68-2.39 points improvement), especially under high pruning ratios, highlighting its importance in preventing conflicting information during recovery training. These improvements are consistently observed across different pruning schemes, with particularly pronounced effects in structured pruning where capability degradation tends to be more severe and uneven.
@@
Recovery Post-training Efficiency Analysis To highlight PASER’s advantages on recovery post-training efficiency, we conduct the experiments under different data budgets $B$ and different datasets and record the corresponding averaged reasoning performance and training time in Figure 2. From the first and third subfigures, we can observe that PASER can obtain best recovery performance under different $B/N$ on Alpaca and LaMini. Interestingly, in the first subfigure, when rising $B/N$ from 0.3 to 0.4, the reasoning performance of Random even decreases. It is because expanding data scale also introduces the conflicting or negative data existing in the original dataset. From the second and fourth subfigures, PASER consistently consumes the least training time, which can be attributed to the efficiency-driven sample selection process in PASER. This advantage can be more obvious under low $B/N$ like 0.02 on LaMini. This is because increasing data budget will force PASER to select some relatively more time-consuming samples given the fixed original dataset, weakening its efficiency superiority.
In addition, more experiment results and analysis can be found in Appendix C, D, E, F, G, H, I, and J.
@@ 5 Conclusion and Future Works
Recovery post-training has been an important procedure after large language model pruning to restore the critical capabilities. Previous works directly utilize the full instruction tuning dataset, facing high computation cost and risks of untargeted recovery and negative transfer. In this work, we propose the post-training data selection method for efficient pruned model recovery. According to capability degradation degrees, we allocate selection budget among different capability data obtained through semantic-structural clustering. We then select samples where model behavior has been severely affected while considering computation cost, and introduce a concept consistency graph to mitigate negative transfer. Extensive experiments on different LLMs demonstrate the effectiveness of our framework. Future work will explore other optimization approaches like data augmentation and revision to further improve recovery efficiency.
Impact Statement
This work aims to improve the efficiency of recovering pruned large language model capabilities while reducing computational costs. Our approach has several broader implications for society and the field of machine learning: On the positive side, PASER could help reduce the environmental impact of training large language models by enabling more efficient recovery processes that require only 4-20% of the original post-training data. This aligns with growing efforts to make AI development more environmentally sustainable. Additionally, by reducing computational requirements, our method could make working with large language models more accessible to researchers and organizations with limited computing resources, potentially democratizing access to AI technology. However, we acknowledge potential risks and challenges. As our method helps recover model capabilities more efficiently, it could accelerate the deployment of compressed language models in real-world applications. This warrants careful consideration of model safety, reliability, and potential misuse. While PASER focuses on maintaining model performance, future work should examine how capability recovery impacts model biases and fairness. We encourage implementing appropriate safeguards when deploying recovered models in practice. We also note that as this technology develops, it will be important to monitor its effects on energy consumption patterns in AI development and establish best practices for responsible use. The AI community should continue discussing how efficiency improvements like PASER can best serve society’s interests while minimizing potential harms. Our hope is that by making language model compression more practical and efficient, this work will contribute to more sustainable and accessible AI development, while acknowledging the need for careful consideration of safety and ethical implications in its deployment.
@@ Appendix A Supporting Materials for Introduction Part
Figure 3: Average performance on seven common LLM reasoning evaluation tasks after recovery post-training with different data. The numbers in brackets represent the group index of the data subset in the full dataset. Unpruned indicates the original model and w/o Training indicates the pruned model (using LLM-Pruner (Ma et al., 2023)) without the recovery post-training.
@@ Appendix B Pseudocode
0:  $M_{o}$: original model, $M_{p}$: pruned model, $D$: instruction tuning dataset, $B$: data budget, $U$: computational cost budget
  function PASER ($M_{o},M_{p},D,B,U$
@@
        $\text{CDS}(c_{k})\leftarrow\text{CapabilityDegradationScore}(c_{k},M_{o},M_{p})$

### v2

@@ 3.4 Negative Tuning Effects Mitigation
A CCG is a graph $G=(V,E)$ where the vertices $V$ represent concepts, and an edge $(v_{i},v_{j})\in E$ exists if concepts $v_{i}$ and $v_{j}$ co-occur in at least one instruction tuning sample without conflict.
For each new sample $(x,y)$, we first extract its concept $C(x,y)$ and then check for consistency: $\text{IsConsistent}(x,y)=\forall v_{i},v_{j}\in C(x,y):(v_{i},v_{j})\in E\text{ or }\{v_{i},v_{j}\}\not\subset V$. We only add samples that are consistent with the existing CCG, updating the graph with each addition. This approach ensures that we maintain a coherent set of instructions, minimizing the risk of negative tuning effects by avoiding the introduction of conflicting concepts during the recovery training process. The full version of our algorithm is provided in Appendix B.
@@ 3.5 Time Complexity Analysis
As the main bottleneck for application, we provide a theoretical analysis for PASER’s time complexity.
@@
Proof. We analyze each step of the algorithm in detail. The Semantic-Structural Recovery Instruction Clustering, including SentenceBERT embedding, Diffusion Kernel computation, and NMF Spectral Clustering, has a dominant complexity of $O(N\log N)$. For the Capability Degradation Assessment step, computing JSD for each sample and calculating CDS for each cluster take $O(N)$ time in total. The Inter-capability Budget Allocation, which involves allocating the budget to clusters, has a time complexity of $O(K)$, where $K$ is the number of clusters. However, since $K\leq N$, this step does not dominate the overall complexity. During Intra-capability Efficiency-Driven Sample Selection, for each cluster $c_{k}$, we perform sorting by JSD ($O(|c_{k}|\log|c_{k}|)$), iterate through sorted samples ($O(|c_{k}|)$), perform consistency checks (IsConsistent, $O(C^{2})$ per sample), and update the CCG ($O(C^{2})$ per sample). Considering all clusters, this step’s total complexity is $O(N\log N+NC^{2})$. Thus, the overall time complexity is dominated by the clustering step and the intra-capability sample selection step. Therefore, the total time complexity is $O(N\log N+NC^{2})$.
In practice, $C$ is often much smaller than $N$ ($C\ll N$) and can be considered as a constant factor for large $N$. Thus, we can simplify the complexity to $O(N\log N)$. This analysis demonstrates that PASER is computationally efficient and scalable for large instruction tuning datasets. Besides, we also conduct a comprehensive theoretical error bound analysis for PASER in Appendix C.
@@ 4.1 Experiment Setup
Target LLMs The experiments are mainly performed on several open-source popular English LLMs: LLaMA2-7B/13B/70B (Touvron et al., 2023) (hf version), LLaMA3-8B (Dubey et al., 2024)(instruct version), and bilingual LLMs: Baichuan2-7B/13B (Yang et al., 2023)(base version), which support both English and Chinese. Besides, several more recent LLMs are also adopted in Appendix H.
@@
Base Pruning Schemes Different pruning schemes are incorporated into our experiments to explore the applicability of PASER, ranging from structured pruning methods: LLM-Pruner (Ma et al., 2023), SliceGPT (Ashkboos et al., 2024), semi-structured pruning method: Wanda (Sun et al., 2024), and unstructured pruning method: SparseGPT (Frantar and Alistarh, 2023).
@@
Evaluation Datasets and Tasks To thoroughly evaluate the performance of recovered LLMs, we employ seven common sense reasoning datasets:BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018). In the practice, we relies on the open-source library [^1] https://github.com/EleutherAI/lm-evaluation-harness to implement the evaluation, where the model needs to rank the choices in the multiple choice tasks or generate the answer in the open-ended generation tasks. The whole process is conducted in the zero-shot manner. Besides, we follow  (Ma et al., 2023) to evaluate the language modeling capability with the zero-shot perplexity (PPL) analysis on WikiText2 (Merity et al., 2022) and PTB (Marcus et al., 1993). For enhancing evaluation comprehensiveness, we also conduct experiments on mathematical problem solving and code generation tasks, as shown in Appendix D and E, respectively.
@@ 4.2 Experiment Results and Analysis
Recovery Performance for Different Pruning Schemes We evaluate the recovery performance of LLaMA2-7B using different instruction tuning data selection methods under structured pruning, semi-structured pruning, and unstructured pruning, respectively. According to the results in Table 1, directly employing full data can indeed bring the sub-optimal recovery performance, especially under the LLM-Pruner. This is because the full version of data contains some irrelevant or conflicting information for capability recovery, resulting in the negative tuning effects during the training phase. Meanwhile, even the general instruction tuning data selection methods like IFD and Nuggets can bring better reasoning and language model performance than full data and random in most cases, validating the necessity of conducting recovery data selection. Furthermore, we can find that previous selection methods can hardly help model recover to the level of unpruned status, under the limited data budget. However, our PASER can not only outperform baselines, but also reduce the averaged reasoning performance degradation to less than 3% under LLM-Pruner, Wanda, and SparseGPT. Especially, when pruning LLaMA2-7B with SliceGPT, our PASER can improve the average reasoning performance to 64.31, higher than the unpruned model. Besides, its zero-shot perplexity on WikiText2 and PTB is also lower than unpruned model slightly. This suggests that allocating recovery budget according to capability degradation levels and prioritizing most-affected samples exhibit the potential of help pruned LLM recover to the capability level of unpruned status. Besides, PASER can also be extended to other LLM post-compression scenarios, like the post-quantization recovery. The corresponding results and analysis are provided in Appendix F.
@@
Robustness over Various Target Large Language Models To validate whether PASER can maintain the robust effectiveness among various target LLMs, we conduct the experiments on LLaMA2-7B/13B/70B, LLaMA3-8B, Baichuan2-7B/13B, under LLM-Pruner. According to results in Table 2, we can first observe the model capability under high pruning ratio (50%) is hard to recover to unpruned level, especially for relatively smaller model like LLaMA2-13B and Baichuan2-13B. Though, PASER can still outperform random and best-performing data selection baseline, Nuggets, by 4.41 and 2.31 points, respectively on average. Especially, for LLaMA2-70B, our PASER can control the averaged reasoning performance degradation to less than 3%. This can be explained that the structure redundancy in 70B model is relatively higher, paving the way for effective recovery through data selection under high pruning ratios. As for the second smallest model, LLaMA3-8B, PASER can recover the reasoning performance to the 99.56% of the unpruned status, which further demonstrates the robustness of PASER over different target LLMs. Finally, the performance of various recovery methods including PASER on Baichuan2-7B is not satisfying enough given only 25% pruning ratio, which can be attributed to that the pruning process has severely damaged the model internal structure. Supplementary experiments on more recent LLMs can be seen in Appendix H.
Table 3: Recovery performance of different instruction tuning data selection methods under two structured pruning schemes on LLaMA2-7B model. The ‘bold’ represents the best performance under the same pruning scheme. Here, the LaMini is taken as the original dataset. The “Reason” indicates the averaged performance on 7 reasoning datasets.
@@
Figure 2: Average reasoning performance and recovery post-training time consumption curves corresponding to different instruction tuning data selection methods. The left two subfigures are for Alpaca while right two subfigures are for LaMini.
@@
Ablation Study To validate the contribution of each component in PASER, we conduct comprehensive ablation studies. Specifically, we evaluate three variants: (1) PASER w/o S2RIC: replacing semantic-structural clustering with random clustering while keeping other modules unchanged; (2) PASER w/o CDAIS: randomly sampling equal number of instructions from each cluster instead of using capability degradation-aware selection; (3) PASER w/o NTEM: removing the negative tuning effects mitigation module. The results in Table 4 demonstrate that all three components contribute positively to model recovery across different pruning schemes. The semantic-structural clustering effectively identifies capability-specific instruction groups, leading to 0.18-1.43 points improvement in reasoning performance. Its removal causes degradation in both language modeling (increased perplexity) and reasoning tasks, particularly evident under structured pruning schemes like LLM-Pruner and SliceGPT. The capability degradation-aware selection mechanism enhances recovery efficiency through adaptive budget allocation, contributing 0.29-1.63 points improvement in reasoning tasks while maintaining stable language modeling performance. Negative tuning effects mitigation shows significant impact (0.68-2.39 points improvement), especially under high pruning ratios, highlighting its importance in preventing conflicting information during recovery training. These improvements are consistently observed across different pruning schemes, with particularly pronounced effects in structured pruning where capability degradation tends to be more severe and uneven. More detailed ablation study results and analysis are provided in Appendix I. Besides, we also empirically explore the technique selection of PASER’s each component in Appendix K.
@@
Recovery Post-training Efficiency Analysis To highlight PASER’s advantages on recovery post-training efficiency, we conduct the experiments under different data budgets $B$ and different datasets and record the corresponding averaged reasoning performance and training time in Figure 2. From the first and third subfigures, we can observe that PASER can obtain best recovery performance under different $B/N$ on Alpaca and LaMini. Interestingly, in the first subfigure, when rising $B/N$ from 0.3 to 0.4, the reasoning performance of Random even decreases. It is because expanding data scale also introduces the conflicting or negative data existing in the original dataset. From the second and fourth subfigures, PASER consistently consumes the least training time, which can be attributed to the efficiency-driven sample selection process in PASER. This advantage can be more obvious under low $B/N$ like 0.02 on LaMini. This is because increasing data budget will force PASER to select some relatively more time-consuming samples given the fixed original dataset, weakening its efficiency superiority. Besides, we also study the time consumption during data selection in Appendix L.
@@ 5 Conclusion and Future Works
Recovery post-training has been an important procedure after large language model pruning to restore the critical capabilities. Previous works directly utilize the full instruction tuning dataset, facing high computation cost, risks of untargeted recovery, and negative tuning effects. In this work, we propose the post-training data selection method for efficient pruned model recovery. According to capability degradation degrees, we allocate selection budget across different capability data obtained through semantic-structural clustering. We then select samples where model output behavior has been severely affected while considering computation cost, and introduce a concept consistency graph to mitigate negative tuning effects. Extensive experiments on different LLMs and theoretical analysis demonstrate the effectiveness of our framework. Future work will explore other optimization approaches like data augmentation and revision to further improve recovery efficiency.
@@ Appendix A Supporting Materials for Introduction Part
Figure 3: Average performance on seven common LLM reasoning evaluation tasks after recovery post-training with different data. The numbers in brackets represent the group index of the data subset in the full dataset. Unpruned indicates the original model and w/o Training indicates the pruned model (using LLM-Pruner [Ma et al., 2023]) without the recovery post-training.
@@ Appendix B Pseudocode
Input: $M_{o}$: original model, $M_{p}$: pruned model, $D$: instruction tuning dataset, $B$: data budget, $U$: computational cost budget
procedure PASER($M_{o},M_{p},D,B,U$)
@@
         $\text{CDS}(c_{k})\leftarrow\text{ComputeCapabilityDegradationScore}(c_{k},M_{o},M_{p})$